You have recently been hired by KEEA Manufacturing to work in its newly established treasury department. KEEA Manufacturing is a small company that produces highly customized cardboard boxes in a variety of sizes for different purchasers. Ibrahim KEEA, the owner of the company, works primarily in the sales and production areas of the company. Currently, the company basically puts all receivables in one pile and all payables in another, and a part-time bookkeeper periodically comes in and works on the piles. Because of this disorganized system, the finance area needs work, and that’s what you’ve been brought in to do. The company currently has a cash balance of GHS170,000, and it plans to purchase new machinery in the third quarter at a cost of GHS300,000. The purchase of the machinery will be made with cash because of the discount offered for a cash purchase. Ibrahim wants to maintain a minimum cash balance of GHS130,000 to guard against unforeseen contingencies. All of KEEA sales to customers and purchases from suppliers are made with credit, and no discounts are offered or taken. The company had the following sales each quarter of the year

Quarter 1 Quarter 2 Quarter 3 Quarter 4

Gross sales GHS937,000 GHS968,000 GHS1,032,000 GHS908,000

After some research and discussions with customers, you’re projecting that sales will be 8 percent higher in each quarter next year. Sales for the first quarter of the following year are also expected to grow at 8 percent. You calculate that KEEA currently has an accounts receivable period of 57 days and an accounts receivable balance of GHS639,000. However, 10 percent of the accounts receivable balance is from a company that has just entered bankruptcy, and it is likely that this portion will never be collected.
You’ve also calculated that KEEA typically orders supplies each quarter in the amount of 50 percent of the next quarter's projected gross sales, and suppliers are paid in 53 days on average. Wages, taxes, and other costs run about 25 percent of gross sales. The company has a quartely interest payment of GHS180,000 on it long-term debt. Finally the company uses a local bank for its short-term financial needs.
In the budgeting years, KEEA estimates that wages and salaries of employees will be GHS50,000 per quarter and the company will also pay rent at GHS 40,000 per quarter.


Required:
Prepare a cash budget for KEEA Manufacturing for the next four quarters.
Answer #1
Computation of Quarterly Cash Receipts
WORKING NOTE 1 - Sales
Sales of Qtr 4 (Last Year) Sales of Qtr 1 (Next Year) Sales of Qtr 2 (Next Year) Sales of Qtr 3 (Next Year) Sales of Qtr 4 (Next Year)
            9,08,000                      9,80,640             10,59,091              11,43,818          12,35,324
(Given) (908000*1.08) 980640*1.08 1059091*1.08 1143818*1.08
(8% Growth) (8% Growth) (8% Growth) (8% Growth)
WORKING NOTE 2 - Collection ( Accounts Receivable)
Collection from accounts receivable 5,75,100 (639000*(1-0.1)
Provision of 10% Considered
WORKING NOTE 3
--> Collection during each quarter being the first months' sale which comes out to be (33/90)of quarterly sales (Considering 57 days collection period)
--> (57/90)of quarterly sales collected in next quarter
Qtr 4(Last Year) Qtr 1 Qtr 2 Qtr 3 Qtr 4 Qtr1 suceeding year
A Sales(GHS) (WN1) 9,08,000 9,80,640 10,59,091 11,43,818 12,35,324 13,34,150
B Collection from last years accounts receivable (WN2) 5,75,100
C=(33/90)*A Collection from current quarter sales (WN3) 359568 388333 419400 452952
D=(57/90)*A(Previous Qtr) Collection from previous quarter (WN3) 6,21,072 6,70,758 7,24,418
E=B+C+D Total Cash Receipts 9,34,668 10,09,405 10,90,158 11,77,370
CALCULATION OF CASH EXPENSES ON SUPPLIES
Qtr 1 Qtr 2 Qtr 3 Qtr 4
F Cost of Supplies (50% of next quarter sales)                      5,29,546               5,71,909                6,17,662            6,67,075
G Cash payment during the quarter (supplies made) 529546 571909 617662 667075
CASH BUDGET
Qtr 1 Qtr 2 Qtr 3 Qtr 4
H Opening Balance                      1,70,000               1,30,000                1,30,000            1,30,000
I Cash Receipts                      9,34,668             10,09,405              10,90,158          11,77,370
J=H+I Total cash available                    11,04,668            11,39,405             12,20,158          13,07,370
Cash Payments:
K Payments for supplies                      5,29,546               5,71,909                6,17,662            6,67,075
L Purchase of machinery                3,00,000
M=0.25*A Wages,taxes and other costs                      2,45,160               2,64,773                2,85,955            3,08,831
N Quarterly interest payment                      1,80,000               1,80,000                1,80,000            1,80,000
P Wages and Salaries                          50,000                  50,000                   50,000                50,000
Q Payments for Rent                          40,000                  40,000                   40,000                40,000
R=K+L+M+N+P+Q Total cash payments                    10,44,706            11,06,682             14,73,617          12,45,906
S=J-R Excess /(shortage ) of Cash                          59,963                  32,724               -2,53,459                61,464
T Minimum Balance Required                      1,30,000               1,30,000                1,30,000            1,30,000
U=T-S Short Term Loan/Borrowing                          70,038                  97,276                3,83,459                68,536
V=S+U Ending Balance                      1,30,000              1,30,000                1,30,000            1,30,000
Please Discuss in case of Doubt
Best of Luck. God Bless
Please Rate Well